OpenStudy (dls):

The area bounded by the curves y=cosx and y= sinx between the ordianates x=0 and x=3pi/2 is?

OpenStudy (dls):

 \[\LARGE \int\limits_0 ^ \frac{\pi}{2}(\cos x-\sin x)dx+\int\limits_\frac{\pi}{4}^\frac{\pi}{2}(\sin x-\cos x)dx\] My attempt^ I guess i'm missing something because getting wrong answer!